Mathematics 133 – Intensive Calculus for Science 1 Quiz 1 – September 9, 2005

Your Name:

Do all work on this sheet (front and back if needed). You have 20 minutes. No calculators on this one.

A) The number N of bacteria in a culture (in 1000’s) is a function of time (in hours), so N = f(t). Explain the meaning of the equation f(4) = 20 in terms of the number of bacteria present.

B) The price of a computer stock was $45 on September 9. That day a flaw was discovered in the company’s most popular computer and the stock price began to decrease. Give a formula for the price t days later, and say what type of function you have if

1) the price decreases by 2.50% per day

2) the price decreases by $2.50 per day